# AI Agentic Programming: A Survey of Techniques, Challenges, and Opportunities

### v1

@@ 2.3. Agency in AI Systems
Classical AI research has explored agency extensively in domains such as planning, robotics, and multi-agent systems (li2024survey, ; wu2024autogen, ). Traditional agent models focus on several key attributes:
  • Reactivity: the ability to respond to changes in the environment,
  • Proactivity: the pursuit of long-term goals,
  • Social ability: the capacity to communicate and coordinate with other agents or humans,
  • Autonomy: the ability to operate without direct human intervention.
In the context of AI agentic programming, agency manifests through an LLM-based system’s ability to:
  • Interpret open-ended tasks described in natural language,
  • Plan sequences of development steps (e.g., writing, testing, debugging),
  • Invoke and coordinate external tools (e.g., compilers, test runners, linters),
  • Adapt actions based on feedback from the environment (e.g., compiler errors or test failures),
  • Maintain coherent state and reasoning across multiple iterations.
Unlike symbolic AI agents that rely on explicitly defined world models and search-based planning, LLM-based coding agents operate in a probabilistic, language-driven manner. Despite this difference, they increasingly exhibit behaviors aligned with classical definitions of agency, especially when augmented with memory, tool-use modules, and planning routines.
2.3.1. Comparison with Robotics and Reinforcement Learning Agents
Robotic agents typically interact with the physical world through sensors and actuators. Their perception, control, and planning modules are tightly coupled with real-time feedback and often require safety guarantees. Reinforcement learning (RL) agents (sun2024llm, ), by contrast, learn behaviors by maximizing cumulative reward through trial and error, often in simulated environments. These agents explore large state-action spaces and learn policies over time.
Agentic programming systems share similarities with both paradigms:
  • Like robotic agents, coding agents must coordinate perception (e.g., task understanding) and action (e.g., code edits, tool invocations) in an environment with constraints and feedback.
  • Like RL agents, agentic systems benefit from feedback loops (e.g., test results, compiler outputs) and may incorporate exploration, retry strategies, or even reward-guided behavior.
However, coding agents operate in a symbolic, tool-rich environment, where actions are language-based, and environments (e.g., codebases, APIs, test harnesses) are highly structured. They must reason not only about immediate feedback but also about abstract software goals, dependencies, and long-term coherence across multiple steps. This makes agency in programming both uniquely challenging and distinctly different from physical-world or simulated agents.
@@ 2.4.1. Large Language Models
Model Size (B) Context Win. Tool use Provider (access) Open Source MoE Used in coding IDEs GPT-5 N/A 1 M OpenAI (API only) VS Code, Cursor, other IDEs GPT-4 variants (o3, o4, etc.) N/A 128k OpenAI (API only) VS Code, JetBrains, Cursor Claude 4 Opus $\sim$300 200k Anthropic (API) Cursor, Replit (chat) Gemini 2.5 Pro $\sim$200 1M Google (API) Replit, Google Colab Grok 4 $\sim$1.7T $\sim$128k xAI Not publicly integrated DeepSeek R1-0528 671 (act. 37) 160k DeepSeek (API + weights) Emacs, VS Code (via extension) Kimi K2 1000 (act. 32) 128k Limited Moonshot AI (API) Custom plugin support Qwen3-235B-A22B 235 (act. 22) 32k Limited Alibaba Alibaba Cloud IDE Qwen3-Coder-480B-A35B-Instruct 480 (act. 35) 256k Alibaba Alibaba Cloud IDE Solar-Pro 72 128k Upstage (weights on HF) VS Code (via third-party) Openhands-LM-32B-v0.1 32 128k OpenHands VS Code (via extension) Devstral-Large 123 128k Mistral (API only) VS Code (via API) Devstral-Small 24 128k Mistral (API only) VS Code (via API)
LLMs trained on massive corpora of code and natural language form the foundation of modern agentic programming systems. These models, as represented by GPT-5 (openai2025gpt5introducing, ), Claude (claude2024, ), DeepSeek (guo2025deepseek, ), and Gemini (gemini2023, ), serve as the core reasoning engines, powering code generation, task planning, debugging, documentation, and natural language interaction. Their ability to understand and execute complex instructions makes them central to the design of agentic workflows.
Modern LLMs can generate syntactically correct and semantically meaningful code, answer development-related queries, and engage in multi-turn conversations with minimal task-specific fine-tuning. Many of these models leverage few-shot, zero-shot, and in-context learning capabilities, allowing them to generalize across programming languages, frameworks, and task domains. This flexibility enables developers to use the same underlying model for a wide range of software engineering tasks, from scaffolding and unit test generation to bug repair and performance tuning. In addition to general-purpose models, some LLMs like Grok (grok4, ) and Calude Opus (anthropic_opus4, ) are increasingly optimized for coding tasks through specialized instruction tuning, extended context length, tool use capabilities, and integration with retrieval-based systems. These enhancements make them suitable for multi-turn reasoning, code synthesis grounded in external context, and tool-augmented workflows.
Table 1 provides a comparative overview of state-of-the-art LLMs widely used in code-related tasks. The models listed include Grok 4, GPT-5 (GPT-4, o3 and o4 variants), Gemini 2.5 Pro, Claude 4 Opus, DeepSeek V3, Kimi K2, and Qwen3. The table compares their key attributes, such as the context window length. As the capabilities of LLMs continue to evolve, selecting and fine-tuning the appropriate foundation model becomes a critical design choice in building reliable, efficient, and adaptive agentic systems.
@@ 2.4.2. Prompt Engineering and Reasoning Strategies
Effective agentic behavior often requires structured prompting techniques to guide LLMs through multi-step reasoning and tool use. Methods such as chain-of-thought prompting, ReAct (reasoning and acting), scratchpad prompting, and modular prompting enable LLMs to plan, reflect, and revise outputs. These techniques allow agents to decompose complex problems, retain intermediate states, and act in ways that are more transparent and controllable.
@@ 2.4.3. Tool Use and API Integration
Tool Type Examples Compiler gcc (GCC, ), clang (Clang, ), javac (Javac, ), tsc (TSC, ) Debugger gdb (GDB, ), lldb (LLDB, ), pdb (PDB, ) Test Framework pytest (Pytest, ), unittest (Unittest, ), Jest (Jest, ), Mocha (Mocha, ) Linter eslint (ESLint, ), flake8 (Flake8, ), black (Black, ), prettier (Prettier, ) Version Control git (Git, ) Build System make (Make, ), cmake (CMake, ), npm (NPM, ), maven (Maven, ) Package Manager pip (Pip, ) yarn (Yarn, ), cargo (Cargo, ) Language Server pyright (Pyright, ), tsserver (TSServer, )
Agentic systems rely heavily on external tools, such as compilers, debuggers, test frameworks, linters, and version control systems, to validate and refine generated code. Integration with these tools via command-line interfaces, language server protocols (LSP), or RESTful APIs enables agents to execute actions, gather feedback, and close the loop between generation and validation. Tool use is a defining feature of agentic programming, supporting iterative workflows and grounding decisions in observable outcomes.
For example, Table 2 lists some of the tools supported by GitHub Copilot Agent. The set of tools ranges from compilation to testing and version control.
@@ 2.4.4. State and Context Management
LLMs operate under fixed context windows, limiting their ability to reason over long histories. Agentic systems therefore incorporate external memory mechanisms to store plans, results, tool outputs, and partial progress. This memory can take the form of vector stores, scratchpads, or structured logs, allowing the agent to recall relevant information across multiple steps and maintain coherence over long-running tasks.
Table 3 compares the context management strategies of mainstream AI coding agents, revealing substantial differences in context size and memory persistence. Tools like GitHub Copilot and Codeium lack persistent memory, instead using transient methods such as sliding windows or dynamic token budgeting. In contrast, agents like SWE-agent, Devika, and OpenDevin employ persistent storage, often via vector databases or structured stores, to support long-term recall of plans, tool outputs, and project history. Some, such as Cursor IDE and Continue.dev, use embedding-based search to retrieve semantically relevant content, while others summarize prior actions to stay within the available context window.
These differences reflect a clear trade-off: smaller context windows typically rely on lightweight retrieval or summarization, whereas larger windows with persistent memory enable richer state tracking but add storage and retrieval overhead.
@@ 2.4.5. Feedback Loops and Self-Improvement
Agentic programming leverages feedback to refine outputs iteratively. Agents may rerun failed tests, revise prompts based on compiler errors, or reflect on past failures to improve future behavior. Some systems incorporate explicit planning, retry mechanisms, or even gradient-based updates through fine-tuning or reinforcement learning. This closed-loop design supports robustness and adaptability in complex programming tasks.
Together, these enablers mark a significant departure from traditional code generation pipelines. They allow LLMs to operate as interactive systems capable of planning, acting, and learning, paving the way for more autonomous and capable software development agents.
@@ 2.5.1. Program Synthesis
Program synthesis has been a foundational approach to automated code generation, traditionally divided into two types: deductive synthesis uses formal specifications to generate provably correct programs, while inductive synthesis learns from input-output examples with symbolic search and logic programming techniques to infer program logic (gulwani2017program, ; jain2022jigsaw, ). Classical synthesis systems like sketching (solar2008program, ) and more recent neural approaches like RobustFill (devlin2017robustfill, ) specialize in generating targeted code snippets that satisfy precise specifications.
However, program synthesis focuses on single-function generation from formal specifications and typically operates in a one-shot generation mode (devlin2017robustfill, ), whereas agentic programming handles multi-step workflows (e.g., planning, tool use, and iterative refinement) and engages in continuous interaction with development environments (bouzenia2024repairagent, ).
@@
Agent Context Window Persistent Memory Context Management Mechanism GitHub Copilot 16k Sliding window over active buffer Codeium 32k Dynamic token budgeting based on file proximity and edit history Cursor IDE 128k Semantic search over project history SWE-agent 16k Vector DB retrieval for tool outputs and plan state Devika 32k Structured memory via SQLite and embeddings AutoDev 16k summarization of prior actions and tool logs Continue.dev 32k Embedding-based local recall over recent edits OpenDevin 32k RAG over command history, plans, and intermediate outputs
@@ 2.5.2. Code Completion Tools
Code completion, as one of the most commercially successful applications of AI in programming, excels at context-aware code suggestion, leveraging large-scale pre-training on code repositories to predict next tokens of partially written code (DBLP:conf/iclr/codegen23, ; DBLP:journals/corr/abs-2203-07814, ; husein2025large, ). Advanced completion tools can suggest entire functions, classes, or small modules based on comments, function signatures, and surrounding context, with tools like GitHub Copilot (github_copilot, ), TabNine (tabnine, ), and Amazon Q Developer (amazon-q-developer, ) achieving widespread adoption.
These code completion tools often operate as reactive assistants that respond to developer input, while agentic programming systems demonstrate proactive behavior and autonomous planning. Furthermore, agentic programming extends beyond code generation to encompass testing, debugging, deployment, and maintenance activities that completion tools typically do not address (yang2024swe, ; bouzenia2024repairagent, ).
@@ 2.5.3. DevOps Automation
DevOps automation focuses on streamlining software delivery pipelines through Infrastructure as Code (IaC), Continuous Integration/Continuous Deployment (CI/CD), and automated testing frameworks (humble2010continuous, ). Tools like Jenkins (jenkins, ), GitLab CI (gitlab_ci, ), and modern platforms like GitHub Actions (github_actions, ) automate repetitive deployment tasks, testing workflows, and infrastructure management.
While both paradigms emphasize automation, DevOps automation primarily handles pre-defined workflows and infrastructure management, whereas agentic programming focuses on adaptive problem-solving and creative solution generation. Additionally, agentic programming can potentially orchestrate and improve DevOps processes themselves, representing a higher-order form of automation (kambala2024intelligent, ).
2.5.4. AutoML and Automated Development
Automated Machine Learning (AutoML) represents a successful paradigm for democratizing AI model development through automation of model selection, hyperparameter tuning, and feature engineering (he2021automl, ). Platforms like Google Cloud AutoML (google_automl, ), Amazon SageMaker Autopilot (sagemaker_autopilot, ), and open-source frameworks like Auto-sklearn automate the traditional machine learning pipeline from data preprocessing to model deployment (feurer2015efficient, ).
However, AutoML focuses on statistical model optimization within well-defined machine learning workflows and operates with structured data and standardized evaluation metrics, while agentic programming tackles more general software development challenges with creative problem-solving and multi-modal reasoning.
@@ 2.5.5. Multi-Agent Systems and Human-AI Collaboration
Traditional multi-agent systems in software development typically involve specialized agent roles working within predefined coordination protocols (jennings2000agent, ). These systems often feature separate agents for requirements analysis, code generation, testing, and documentation, coordinating through structured communication interfaces. Recent advances in LLM-based multi-agent programming have demonstrated the potential for more sophisticated collaboration, with systems like MetaGPT showing how multiple AI agents can simulate software development teams (hong2023metagpt, ).
AI agentic programming can be viewed as an evolution of multi-agent systems that incorporates human-in-the-loop collaboration and dynamic role adaptation. Unlike traditional multi-agent systems with fixed agent roles and rigid communication protocols, agentic programming systems demonstrate fluid role assignment and context-adaptive behavior. Moreover, the integration of tool use, environmental interaction, and persistent memory distinguishes modern agentic programming from earlier multi-agent approaches. Contemporary agentic systems like AutoGen (wu2024autogen, ) and CrewAI (duan2024exploration, ) enable agents to directly interact with development tools, maintain context across extended sessions, and learn from past interactions (li2024survey, ). This represents a significant advancement over traditional multi-agent systems that typically operated in more constrained, simulation-based environments.

### v2

@@ 2.3. Agency in AI Systems
Classical AI research has explored agency extensively in domains such as planning, robotics, and multi-agent systems (Li et al., 2024e; Wu et al., 2024a). Traditional agent models are typically characterized by four key attributes: reactivity, the ability to respond to changes in the environment; proactivity, the pursuit of long-term goals; social ability, the capacity to communicate and coordinate with other agents or humans; and autonomy, the ability to operate without direct human intervention. In the context of AI agentic programming, these notions of agency are realized in new ways. An LLM-based system can interpret open-ended tasks expressed in natural language, plan sequences of development steps such as writing, testing, and debugging, and invoke or coordinate external tools like compilers, test runners, and linters. It can also adapt its actions in response to environmental feedback, such as compiler errors or test failures, while maintaining coherent state and reasoning across multiple iterations. Unlike symbolic AI agents that rely on explicitly defined world models and search-based planning, LLM-based coding agents operate in a probabilistic, language-driven manner. Despite this difference, they increasingly exhibit behaviors aligned with classical definitions of agency, especially when augmented with memory, tool-use modules, and planning routines.
@@ 2.4.1. Large language models
[…]
LLMs trained on massive corpora of code and natural language form the foundation of modern agentic programming systems. These models, as represented by GPT-5 (OpenAI, 2025), Claude (Anthropic, 2024b), DeepSeek (Guo et al., 2025), and Gemini (Anil et al., 2023a), serve as the core reasoning engines, powering code generation, task planning, debugging, documentation, and natural language interaction. Their ability to understand and execute complex instructions makes them central to the design of agentic workflows. Modern LLMs can generate syntactically correct and semantically meaningful code, answer development-related queries, and engage in multi-turn conversations with minimal task-specific fine-tuning. Many of these models leverage few-shot, zero-shot, and in-context learning capabilities, allowing them to generalize across programming languages, frameworks, and task domains. This flexibility enables developers to use the same underlying model for a wide range of software engineering tasks, from scaffolding and unit test generation to bug repair and performance tuning. In addition to general-purpose models, some LLMs like Grok (xAI, 2025) and Calude Opus (Anthropic, 2025b) are increasingly optimized for coding tasks through specialized instruction tuning, extended context length, tool use capabilities, and integration with retrieval-based systems. These enhancements make them suitable for multi-turn reasoning, code synthesis grounded in external context, and tool-augmented workflows. Table 1 provides a comparative overview of some of the state-of-the-art LLMs used in code-related tasks. The table compares their key attributes. As the capabilities of LLMs continue to evolve, selecting and fine-tuning the appropriate foundation model for coding tasks becomes a critical design choice in building reliable, efficient, and adaptive agentic systems.
@@ 2.4.2. Prompt engineering and reasoning strategies
Effective agentic behavior often requires structured prompting techniques to guide LLMs through multi-step reasoning and tool use. Rather than relying on a single input-output exchange, these methods provide scaffolding that helps models break tasks into manageable steps and maintain coherence across longer interactions. For instance, chain of thought prompting (Wei et al., 2022) encourages explicit reasoning traces, making intermediate steps visible and improving problem-solving accuracy. ReAct (reasoning and acting) interleaves reasoning with concrete actions (Yao et al., 2023c), such as tool calls or environment interactions, enabling agents to both deliberate and act in context. Scratchpad prompting (Abbe et al., 2024) provides a working memory space where partial results, hypotheses, or plans can be written down and refined, supporting iterative refinement. Modular prompting (Sun et al., 2023a), on the other hand, separates tasks into distinct functional roles—such as planner, executor, and verifier—so that the model can coordinate across specialized subtasks. Together, these techniques allow agents to decompose complex problems, retain intermediate states, and revise their behavior in light of new evidence. They also increase transparency by exposing the reasoning process and provide greater controllability by constraining how models structure their outputs. In practice, structured prompting forms the backbone of many agentic systems, enabling LLMs to move beyond ad hoc responses and toward more reliable, interpretable, and goal-directed behavior.
@@ 2.4.3. Tool use and API integration
[…]
import OpenAI from "openai";
const client = new OpenAI();
const tools = [
"type": "function",
"function": {
"name": "compile_code",
"description": "...",
"parameters": {
"type": "object",
"properties": {
"language": {
"type": "string",
"enum": ["c", "cpp"],
"description": "Programming language to compile."
"source": {
"type": "string",
"description": "Source code to compile."
"flags": {
"type": "array",
"items": { "type": "string" },
"description": "Compiler flags."
"required": ["language", "source"],
"additionalProperties": false
Listing 1: An example of the OpenAI tool schema
Table 3. Example interfaces between LLMs and tools.
[…]
Agentic systems rely heavily on external tools, such as compilers, debuggers, test frameworks, linters, and version control systems, to validate and refine generated code. These tools provide the concrete signals needed to check correctness, enforce coding standards, and ensure that outputs remain consistent with project requirements. For example, Table 2 lists a subset of the tools currently supported by the GitHub Copilot Agent (GitHub, 2021), covering compilation, testing, and version control. Integration with external tools can take multiple forms, including command line interfaces, language server protocols (LSP), and RESTful APIs. Increasingly, LLMs interact with tools through structured Python or JavaScript interfaces that specify the available actions, input parameters, and expected outputs in a machine-readable format. For example, Listing 1 shows how a compiler can be exposed as a tool to extend an LLM’s capabilities. This structured approach reduces ambiguity, grounds commands in the correct syntax, and makes it easier for the model to call external tools safely and consistently. By interpreting the schema, an LLM can generate well-formed commands, parse structured responses, and adjust its behavior in a predictable way. Table 3 summarizes the main types of interfaces through which LLMs interact with external tools. These range from free-form natural language instructions to highly structured schemas and domain-specific protocols.
@@ 2.4.4. State and context management
LLMs operate under fixed context windows, limiting their ability to reason over long histories. Agentic systems therefore incorporate external memory mechanisms to store plans, results, tool outputs, and partial progress. This memory can take the form of vector stores, scratchpads, or structured logs, allowing the agent to recall relevant information across multiple steps and maintain coherence over long-running tasks. Table 4 compares the context management strategies of mainstream AI coding agents, revealing substantial differences in context size and memory persistence. Tools like GitHub Copilot (GitHub, 2021) currently do not utilize persistent memory, instead using transient methods such as sliding windows or dynamic token budgeting. In contrast, agents like SWE-agent (Yang et al., 2024b), Devika (dev, 2025), and OpenDevin (Ope, 2025) employ persistent storage, often via vector databases or structured stores, to support long-term recall of plans, tool outputs, and project history. Some, such as Cursor IDE (Cursor, 2023) and Continue.dev (con, 2025), use embedding-based search to retrieve semantically relevant content, while others summarize prior actions to stay within the available context window. These differences reflect a clear trade-off: smaller context windows typically rely on lightweight retrieval or summarization, whereas larger windows with persistent memory enable richer state tracking but add storage and retrieval overhead.
@@ 2.4.5. Feedback loops and self-improvement
Agentic programming leverages feedback to refine outputs iteratively. Agents may rerun failed tests, revise prompts based on compiler errors, or reflect on past failures to improve future behavior. For instance, compiler errors may trigger targeted code edits, test failures may prompt iterative debugging, and linter warnings may guide stylistic refinements. Some systems incorporate explicit planning, retry mechanisms, or even gradient-based updates (Huang et al., 2024c) through fine-tuning or reinforcement learning. This closed-loop design supports robustness and adaptability in complex programming tasks.
@@ 2.5.1. Program synthesis
Program synthesis has been a foundational approach to automated code generation, traditionally divided into two types: deductive synthesis uses formal specifications to generate provably correct programs, while inductive synthesis learns from input-output examples with symbolic search and logic programming techniques to infer program logic (Gulwani et al., 2017; Jain et al., 2022). Classical synthesis systems like sketching (Solar-Lezama, 2008) and more recent neural approaches like RobustFill (Devlin et al., 2017) specialize in generating targeted code snippets that satisfy precise specifications. However, classical program synthesis focuses on single-function generation from formal specifications (due to the scaling challenge of code sysnthesis (Gulwani et al., 2017)) and typically operates in a one-shot generation mode (Devlin et al., 2017), whereas agentic programming handles multi-step workflows (e.g., planning, tool use, and iterative refinement) and engages in continuous interaction with development environments (Bouzenia et al., 2024).
@@
[…]
@@ 2.5.2. Code completion tools
Code completion, as one of the most commercially successful applications of AI in programming, excels at context-aware code suggestion, leveraging large-scale pre-training on code repositories to predict next tokens of partially written code (Nijkamp et al., 2023; Li et al., 2022; Husein et al., 2025). Advanced completion tools can suggest entire functions, classes, or small modules based on comments, function signatures, and surrounding context, with tools like GitHub Copilot (GitHub, 2021), TabNine (Tabnine, 2025), and Amazon Q Developer (Amazon Web Services, 2025) achieving widespread adoption. These code completion tools often operate as reactive assistants that respond to developer input, while agentic programming systems demonstrate proactive behavior and autonomous planning. Furthermore, agentic programming extends beyond code generation to encompass testing, debugging, deployment, and maintenance activities that completion tools typically do not address (Yang et al., 2024b; Bouzenia et al., 2024).
@@ 2.5.3. DevOps automation
DevOps automation focuses on streamlining software delivery pipelines through Infrastructure as Code (IaC), Continuous Integration/Continuous Deployment (CI/CD), and automated testing frameworks (Humble and Farley, 2010). Tools like Jenkins (Project, 2025), GitLab CI (Inc., 2025), and modern platforms like GitHub Actions (GitHub, 2025) automate repetitive deployment tasks, testing workflows, and infrastructure management. While both paradigms emphasize automation, DevOps automation primarily handles pre-defined workflows and infrastructure management, whereas agentic programming focuses on adaptive problem-solving and creative solution generation. Additionally, agentic programming can potentially orchestrate and improve DevOps processes themselves, representing a higher-order form of automation (KAMBALA, 2024).
2.5.4. Automated machine learning and automated development
Automated Machine Learning (AutoML) represents a successful paradigm for democratizing AI model development through automation of model selection, hyperparameter tuning, and feature engineering (He et al., 2021). Platforms like Google Cloud AutoML (Cloud, 2025a), Amazon SageMaker Autopilot (Services, 2025), and open-source frameworks like Auto-sklearn automate the traditional machine learning pipeline from data preprocessing to model deployment (Feurer et al., 2015). However, AutoML focuses on statistical model optimization within well-defined machine learning workflows and operates with structured data and standardized evaluation metrics, while agentic programming tackles more general software development challenges with creative problem-solving and multi-modal reasoning.
@@ 2.5.5. Multi-Agent systems and human-AI collaboration
Traditional multi-agent systems in software development typically involve specialized agent roles working within predefined coordination protocols (Jennings, 2000). These systems often feature separate agents for requirements analysis, code generation, testing, and documentation, coordinating through structured communication interfaces. Recent advances in LLM-based multi-agent programming have demonstrated the potential for more sophisticated collaboration, with systems like MetaGPT showing how multiple AI agents can simulate software development teams (Hong et al., 2023). AI agentic programming can be viewed as an evolution of multi-agent systems that incorporates human-in-the-loop collaboration and dynamic role adaptation. Unlike traditional multi-agent systems with fixed agent roles and rigid communication protocols, agentic programming systems demonstrate fluid role assignment and context-adaptive behavior. Moreover, the integration of tool use, environmental interaction, and persistent memory distinguishes modern agentic programming from earlier multi-agent approaches. Contemporary agentic systems like AutoGen (Wu et al., 2024a) and CrewAI (Duan and Wang, 2024) enable agents to directly interact with development tools, maintain context across extended sessions, and learn from past interactions (Li et al., 2024e). This represents a significant advancement over traditional multi-agent systems that typically operated in more constrained, simulation-based environments.
2.5.6. Comparison with robotics and reinforcement learning agents
Robotic agents typically interact with the physical world through sensors and actuators. Their perception, control, and planning modules are tightly coupled with real-time feedback and often require safety guarantees. Reinforcement learning (RL) agents (Sun et al., 2024c), by contrast, learn behaviors by maximizing cumulative reward through trial and error, often in simulated environments. These agents explore large state-action spaces and acquire policies over time. Agentic programming systems share features with both paradigms. Like robotic agents, coding agents must coordinate perception, such as task understanding, with actions such as code edits or tool invocations, all within an environment shaped by constraints and feedback. Like RL agents, they benefit from feedback loops, for example, test results or compiler outputs, and may employ exploration, retry strategies, or even reward-guided behavior. At the same time, coding agents differ in multiple ways. They operate in a symbolic, tool-rich environment where actions are language-based and environments, such as codebases, APIs, and test harnesses, are highly structured. Success requires reasoning not only about immediate feedback but also about abstract software goals, dependencies, and long-term coherence across multiple steps. This makes agency in programming uniquely challenging and distinct from both the physical world and simulated agents.